value: 0.00 points Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below Standard Quantity or Standard Price Standard Hours or Rate Direct materials 60 pounds Direct labor 0.4 hours $9.00 per hour Cost 3.00 per pound $18.00 s 3 60 During the most recent month, the following activity was recorded a Eleven thousand six hundred pounds of material were purchased at a cost of $2.90 per pound b. The company produced only 1,160 units, using 10,440 pounds of material. (The rest of the material purchased remained in raw materials inventory) Five hundred and sixty four hours of direct labor time were recorded at a total labor cost of $6,768 c. Required Compute the materials price and quantity variances for the month (Do not round intermediate calculations. Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance).) Direct materials price variance Direct materials quantity varance 1 Stand rd Cost Vorian Type here to search
Answer #1

Material price variance = (3-2.90)*11600 = 1160 F

Material quantity variance = (1160*6-10440)*3 = 10440 U